OpenStudy (anonymous):

Part 1: Using complete sentences, explain the difference between an arithmetic sequence and an arithmetic series. Part 2: Describe a real world scenario that is an arithmetic series. Show your work as you find the 7th term of your scenario.

OpenStudy (anonymous):

Part 1: An arithmetic sequence is a list of numbers which follow a rule. A series is the sum of a sequence of numbers.

OpenStudy (anonymous):

one is a sum the other is not. THe one with the word series is the sum one.

OpenStudy (anonymous):

Part 2: A coffee shop in a city has usually has 10 visitors between 7:00 a.m. and 8:00 a.m. After a while, its customers kept increasing by 2 every day for 7 days in a row. Assume that each time a customer orders one cup of coffee. How many cups of coffee were sold on the seventh day?

OpenStudy (anonymous):

This is an arithmetic series with first term a=10 and common difference d=2. 7th term = a+(7-1)d =a+6d =10+(6)(2) = 22

OpenStudy (anonymous):

lets say i making a pyramid and the the row below has 1 more than the one on top  1 2 3 4 5 that is a sequence the series will be the sum of the numbers in the sequence 1+2+3+4+5